OpenStudy (anonymous):

X^2+10x+21/x^2+8x+15

hero (hero):

\[\frac{x^2+10x + 21}{x^2 + 8x + 15}\]

OpenStudy (anonymous):

yea

hero (hero):

\[\frac{x^2 + 7x + 3x + 21}{x^2 + 3x + 5x + 15}\]

hero (hero):

\[\frac{x(x+7)+3(x+7)}{x(x+3)+5(x+3)}\]

OpenStudy (anonymous):

X^2+10x+21/x^2+8x+15 x^2 +10x +21 = (x+7) (x+3) x^2 +8x + 15 = (x+3)(x+5) X^2+10x+21/x^2+8x+15 = (x+7)(x+3)/ (x+3)(x+5) = (x+7)/(x+5)

hero (hero):

\[\frac{(x+7)(x+3)}{(x+5)(x+3)}\]

OpenStudy (anonymous):

(3 (6 x^3+5 x^2+7))/x^2+X^2

hero (hero):

x+3's cancel

hero (hero):

Final Answer: \[\frac{x+7}{x+5}\]

OpenStudy (anonymous):